Calculate the standard entropy change for the following reaction:

Ca(OH)2(s) ⟶ CaO(s) + H2O(l)



1
Expert's answer
2021-07-08T04:54:30-0400

"ΔS^0 = S^0(CaO) + S^0(H_2O) -S^0(Ca(OH)_2) \\

= 38.2+ 69.91 -83.39 = 24.72 \; J/mol \cdot K"